# HoneySat: A Network-based Satellite Honeypot Framework

### v1

@@ Abstract
Satellites are the backbone of several mission-critical services, such as GPS that enable our modern society to function. For many years, satellites were assumed to be secure because of their indecipherable architectures and the reliance on security by obscurity. However, technological advancements have made these assumptions obsolete, paving the way for potential attacks, and sparking a renewed interest in satellite security. Unfortunately, to this day, there is no efficient way to collect data on adversarial techniques for satellites, which severely hurts the generation of security intelligence.
@@
In this paper, we present HoneySat, the first high-interaction satellite honeypot framework, which is fully capable of convincingly simulating a real-world CubeSat, a type of Small Satellite (SmallSat) widely used in practice. To provide evidence of the effectiveness of HoneySat, we surveyed experienced SmallSat operators currently in charge of active in-orbit satellite missions.
Results revealed that the majority of satellite operators (71.4%) agreed that HoneySat provides realistic and engaging simulations of CubeSat missions. Further experimental evaluations also showed that HoneySat provides adversaries with extensive interaction opportunities by supporting the majority of adversarial techniques (86.8%) and tactics (100%) that target satellites. Additionally, we also obtained a series of real interactions from actual adversaries by deploying HoneySat on the internet over several months, confirming that HoneySat can operate covertly and efficiently while collecting highly valuable interaction data.
1 Introduction
@@
Artificial satellites are complex devices designed to withstand outer space conditions. They serve multiple purposes or types of missions that include position and navigation, e.g., the Global Positioning System (GPS) constellation [74], Earth observation, e.g., the Sentinel constellation [19], and broadband internet service, e.g., the Starlink constellation. In addition, Spacecraft can range from being as massive as thousands of kilograms, such as the International Space Station (ISS) to one kilogram CubeSats. As a result, the software and hardware components that make up a specific spacecraft vary greatly. Likewise, satellite missions’ vary a lot depending on the owner’s budget. For example, university missions are smaller as they have a limited budget. A cyberattack on a satellite or satellite constellation (group of satellites) could have disastrous consequences on a global scale, which is difficult to comprehend. Such an attack could lead to the cessation of air traffic and widespread communication blackouts. It could also cause food shortages and the freezing of financial transactions [11]. Furthermore, such an attack could exacerbate the Kessler Syndrome [18], a scenario in which collisions between satellites and debris in orbit create a cascade effect, generating even more debris, jeopardizing future satellite launches and operations.
@@
Although satellite security research has gained increased attention [81, 60, 53, 79], the relationship between outer space and cyberspace remains poorly understood [52]. At the same time, space-focused threat intelligence remains sparse, and our current methods for identifying tactics, techniques, and procedures (TTPs) used against these critical systems are limited. MITRE ATT&CK currently tracks 152 threat groups but shows only one that targets satellites explicitly [71], highlighting a significant data gap. Although the volume of reported cyber incidents in the space sector has grown, these reports rarely provide sufficient detail [11]. As a result, the security community has limited visibility into adversarial activity aimed at space infrastructure.
Honeypots’ ability to collect real-world cyberattack data makes them an ideal solution to this problem [29]. A honeypot is a decoy computer system intended to lure and entice malicious actors to interact with it [12]; all the while, the honeypot logs all the interactions the attackers make. This log data can later be analyzed to discover new and existing TTPs.
@@
In this paper, we design, implement, and test the first satellite honeypot, HoneySat to attract and analyze adversaries who attack space infrastructures over the Internet, a commonly observed threat vector [7, 5, 33, 80]. HoneySat is a modular, high-interaction honeypot framework that realistically simulates a complete satellite system (ground infrastructure and Satellite). Specifically, HoneySat simulates Small Satellites or SmallSats which are spacecraft with a mass of less than 180 kilograms[44]. CubeSats for example, are SmallSats. Additionally, as part of HoneySat, we developed the Satellite Simulator, a Python project to provide generic simulation functionality for users to populate satellite honeypots with believable data. This capability enables the creation of complete testing environments for satellite software that integrate a variety of simulated subsystems and sensors. For instance, Satellite Simulator can simulate orbital mechanics (e.g., position tracking, attitude adjustment) and electrical power systems (e.g., power generation, consumption, and distribution) alongside other subsystems.
We leveraged our framework to create honeypots of real-world CubeSat missions. Our results, backed by our survey of satellite operators, show that HoneySat’s simulation is highly realistic. Our framework is able to simulate an entire real satellite mission, provides realistic telemetry, and supports real telecommands.
@@
  • A novel framework, HoneySat, a high-interaction, extensible honeypot for small satellites (Sec. 4).
  • The Satellite Simulator, that simulates the physical processes, sensors, and subsystems necessary for a realistic satellite honeypot (Sec. 5).
  • The results of a survey of experienced satellite operators that provide valuable insights into how realistically our honeypot performs, as well as experimental evidence demonstrating that the HoneySat framework can simulate small satellites, collect rich real-world interaction data, and be customized for multiple satellites (Sec. 6).
2 Background
This section lays out key background concepts that are relevant to satellite honeypots. For honeypots: a description of the different existing types (Sec. 2.1), as well as the current state-of-the-art (Sec. 2.2). For satellites: their operation (Sec. 2.3), their architecture (Sec. 2.4), existing Protocol Ecosystems (Sec. 2.5), and tactics, techniques, and procedures (TTPs, Sec. 2.6).
2.1 Types of Honeypots
@@
Low-Interaction Honeypots. These honeypots offer minimal interaction, simulating real systems through scripts or finite-state machines. Their advantages are ease of setup and maintenance due to low resource consumption, and a reduced risk of adversarial takeover. However, they provide limited interaction opportunities to adversaries which limits the interaction data they provide. Low-interaction honeypot examples include Conpot [78] and Honeyd [54].
High-Interaction Honeypots. These honeypots offer extensive interaction opportunities via emulation or advanced simulations [54]. Their main advantage is providing adversaries with almost limitless interactions, enabling them to provide extensive interaction data. However, they pose a high risk of adversarial takeover as adversaries have more opportunities to hijack the honeypot [39]. High-interaction honeypot examples include Cowrie [50] and HoneyPLC [37].
2.2 Honeypot’s State of the Art
The literature on honeypots includes hundreds of implementations that simulate a diverse set of computer systems [30, 48]. From classic implementations that simulate a host’s TCP/IP stack, such as Honeyd [54], to modern approaches that integrate social media applications, such as HoneyTweet [3]. However, there is no satellite honeypot in the literature. In the absence of a satellite honeypot, we now examine the honeypot approaches most related to satellites: Industrial Control Systems (ICS) [36, 35] and Unmanned Aerial Vehicles (UAV), a.k.a., drones summarized in Table 1.
Satellite systems like ICS must be aware of some physical process, e.g., its position, the sun’s position, etc, via sensors to acquire data about the physical world. Several ICS honeypots have simulated these physical processes. For example, ICSnet [58] and HoneyICS [38] simulate several components such as programmable logic controllers (PLCs), and actuators such as water valves.
HoneyDrone [16] is a UAV honeypot that integrates different simulations, e.g., Ardupilot, to recreate multiple attack scenarios, including the UAV ground control station. Although UAV honeypots share some similarities with satellites [75], satellite honeypots require additional physical simulations and involve a more complex operation environment, which we discuss in Sec. 2.3.
Table 1: Comparison of Existing Honeypots and HoneySat.
Keys: = Supported; $\times$× = Not Supported.
@@ 2.2 Honeypot’s State of the Art
4.4,
4.3
4.3,
5.2.1
4.4,
5.1
6.5
2.3 Anatomy of a Satellite Mission
@@
We now describe the components of a satellite mission. Due to satellite missions’ complexity, we explain each component and match it with one of the numbers in Fig. 1. Every satellite mission includes the ground segment from which satellite operators control the satellite and the space segment which includes the satellite itself.
Ground Segment 1. The Ground Segment (GS) covers the terrestrial supporting infrastructure required for a successful satellite operation. It consists of a ground station, responsible for exchanging data with the spacecraft, the computational and network infrastructure required for communication, but also the systems to operate the satellites, e.g., servers, databases, and user interfaces [81]. Several ground stations can be connected in a network and coordinated as part of one GS. The GS includes the Ground Segment Software (GSS) that helps operators schedule and send commands and visualize data that is sent back in response.
Telecommands (TC) 3 and Telemetry (TM) 4. The basic data flow between the space and ground segments are TC and TM [69]. TM is the data the satellite sends to the ground station which may contain the satellite’s status, responses to previous commands, or payload data [69]. TCs are used to operate the satellite and are transmitted and encapsulated in a space protocol (see Sec. 2.5). The design and implementation of TCs varies depending on the satellite mission. From a security perspective, TCs are particularly important as an attacker that can send valid TCs to a satellite can fully take over the mission [81].
Orbital Pass 5. A satellite and its ground station can communicate only during an orbital pass. An orbital pass, or more commonly a pass, is when the satellite rises above a ground station’s horizon and becomes available for communication. A pass’s duration and timing depend on the satellite’s orbit characteristics and any obstructing objects between the satellite and the ground station, e.g., mountains [82]. In some cases, such as geostationary orbit (GEO), the satellite can be above the horizon constantly. The occurrence of passes can be calculated using the spacecraft’s orbital elements. A common representation of the orbital elements is the two-line element set (TLE) [77].
Satellite Mission Operations 6. Satellite mission operations vary widely depending on the owning organization, budget, and technology. Nevertheless, they share some commonalities, which we now describe.
A mission’s operation involves a team of operators that use GSS to operate a satellite or a satellite constellation and ensure the mission’s success [46]. Operations are carried out in a Mission Operations Center (MOC), where satellite operators sit at their workstations to manage TCs sent to the satellite(s).
Satellite operators may also remotely operate satellites 7 by connecting to the ground segment using tools such as desktop sharing like VNC (Virtual Network Computing) [10], or even rely on autonomous or semi-autonomous operations [26].
Satellite operations include two main activities: satellite tracking and TC generation and scheduling. Satellite tracking calculates the satellite’s position in orbit and controls the ground station(s)’ tracking antenna to establish communication between the ground and space segments. TC generation and scheduling crafts command to be programmed and sent to the satellite so it can perform different functions, e.g., read systems status data, schedule orbital maneuvers, schedule payload operations, and download payload data.
@@
Ground (Segment) Software (GSS). GSS allows satellite operators to carry out the satellite mission’s routine operations as described in the previous paragraph. GSS are very diverse. Some satellite missions develop their own GSS while others use open-source [65] or buy proprietary GSS like GSWeb [23]. There are two main types of GSS: Mission Control Software (MCS) and Ground Station Control Software (GSCS).
@@
Mission control software 8 manages TCs and scripts to be sent to the satellite and can display TM. For example, ESA’s SCOS-2000 is an MCS that provides generic functionality that can be customized for a specific mission to cover the functions required for TM reception and processing and TC verification [42]. Some satellite missions develop their own MCS. For example, the SUCHAI mission team developed their MCS application to send and receive TC/TM [65].
@@
Ground station control software (GSCS) 9 helps satellite operators track and visualize the satellite’s orbit and provide detailed information about each satellite’s pass. For example, Gpredict is a popular open-source GSCS that performs real-time satellite tracking and orbit prediction [15]. Gpredict can also interface with and control the radio system and antenna rotor during the satellite pass.
2.4 Satellite Architecture
@@
Satellite architectures are varied and complex, however, here we describe the most common terminology depicted in Fig. 1’s space segment 2.
@@
When referring to satellite architecture there is a distinction between the platform that facilitates the successful operation of the satellite’s day to day-to-day activities and the payload. The platform provides the possibility to run a payload that fulfills the purpose of these missions. This payload differs based on the goal of the mission and can range from instruments to measure physical properties to communications systems providing wireless connectivity.
@@
Platform: The platform, or satellite bus, is a system composed of custom-designed or off-the-shelf subsystems necessary for critical satellite operations. These include, among others, the Attitude Determination and Control System (ADCS) to maintain the satellite’s orientation (i.e. attitude) and position to keep the satellite pointed towards antennas or solar panels illuminated; the Electrical Power Subsystem (EPS) for managing power generation and distribution; the Communication Subsystem (COMM) and the Command and Data Handling (C&DH) subsystems to facilitate communications for receiving TC and sending TM and controlling the satellite operations.
@@
All of these subsystems are then controlled via TCs sent to the satellite. TCs may be, depending on the protocol ecosystem and complexity of the subsystem, interpreted by a central C&DH System or merely forwarded to the recipient subsystem for further processing [84]. This managerial duty is left to Flight Software (FS) running on an embedded system of differing complexity. Some examples include NASA’s Core Flight System (cFS) [41] and F’ (F Prime) [6], KubOS [83], the German Aerospace Center (DLR)’s OUTPOST [17], and the University of Chile’s SUCHAI FS [24]. Attackers aim to gain the ability to send Telecommands to the Flight Software, as this typically grants control over all spacecraft subsystems.
@@
Payload. The payload is the equipment that the satellite employs to fulfill its mission. Due to satellites’ varied missions, payloads are heavily customized [70]. For example, if a satellite’s mission is remote sensing, its payload may include an infrared camera [45]. craft’s essential functions, acts on TCs, and handles internal data transmitted to and from other systems [84]. The CDHS is comprised of Flight Software (FS) running on an OBC [81].
2.5 Small Satellite Protocol Ecosystems

### v2

@@ Abstract
Satellites are the backbone of several mission-critical services that enable our modern society to function, for example, GPS. For years, satellites were assumed to be secure because of their indecipherable architectures and the reliance on security by obscurity. However, technological advancements have made these assumptions obsolete, paving the way for potential attacks, and sparking interest in satellite security. Unfortunately, to this day, there is no efficient way to collect data on adversarial techniques for satellites, hurting the generation of security intelligence that can lead to the development of effective countermeasures.
@@
In this paper, we present HoneySat, the first high-interaction satellite honeypot framework, fully capable of convincingly simulating a real-world CubeSat, a type of Small Satellite (SmallSat). To provide evidence of HoneySat’s effectiveness, we surveyed experienced SmallSat operators in charge of in-orbit satellites and deployed HoneySat over the Internet to entice adversaries.
Our results show that 90% of satellite operators agreed that HoneySat provides a realistic and engaging simulation of a SmallSat mission. Additionally, HoneySat successfully deceived human adversaries in the wild and collected 22 real-world satellite-specific adversarial interactions. Finally, in a major demonstration of HoneySat’s robustness, we collaborated with an aerospace company to perform a hardware-in-the-loop operation that resulted in HoneySat successfully communicating with an in-orbit, operational SmallSat mission.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2026.240537 www.ndss-symposium.org
I Introduction
@@
Satellites are complex devices designed to withstand outer space conditions. They serve multiple purposes or types of missions that include position and navigation, e.g., the Global Positioning System (GPS) constellation [1], Earth observation, e.g., the Sentinel constellation [2], and broadband internet service, e.g., the Starlink constellation [3]. In addition, spacecraft can range from being as massive as thousands of kilograms, such as the International Space Station (ISS), to much smaller one-kilogram CubeSats [4]. As a result, the software and hardware components that make up a specific spacecraft vary greatly. Likewise, satellite missions vary a lot depending on the owner’s budget. For example, university missions are smaller as they have a limited budget [5]. A cyberattack on a satellite or satellite constellation (group of satellites) could have disastrous consequences on a global scale, which is difficult to comprehend. Such an attack could lead to the cessation of air traffic and widespread communication blackouts [6]. It could also cause food shortages and the freezing of financial transactions [7]. Furthermore, such an attack could exacerbate the Kessler Syndrome [8], a scenario in which collisions between satellites and debris in orbit create a cascade effect, generating even more debris, jeopardizing future satellite launches and operations.
@@
Although satellite security research has gained increased attention [12, 13], the relationship between outer space and cyberspace remains poorly understood [14]. At the same time, space-focused threat intelligence remains sparse, and our current methods for identifying the tactics, techniques, and procedures (TTPs) used against these critical systems are limited. MITRE ATT&CK currently tracks 152 threat groups but shows only one that targets satellites explicitly [15], highlighting a significant data gap. Although the volume of reported cyber incidents in the space sector has grown, these reports rarely provide sufficient detail [7]. As a result, the security community has limited visibility into adversarial activity aimed at space infrastructure.
A honeypot’s ability to collect real-world cyberattack data makes it an ideal solution to this problem [16]. A honeypot is a decoy computer system intended to lure and entice malicious actors to interact with it [17]; all the while, the honeypot logs all the interactions the attackers make. This log data can later be analyzed to discover new and existing TTPs.
@@
In this paper, we design, implement, and evaluate the first satellite honeypot, HoneySat, to attract and analyze adversaries who attack space infrastructures over the Internet, a commonly observed threat vector [25, 26, 27, 28]. HoneySat is a modular, high-interaction honeypot framework that realistically simulates a complete satellite system (ground infrastructure and satellite). Specifically, HoneySat simulates Small Satellites or SmallSats, which are spacecraft with a mass of less than 180 kilograms [29]. CubeSats, for example, are SmallSats. Additionally, as part of HoneySat, we developed the Satellite Simulator, a Python project to provide generic simulation functionality for users to populate satellite honeypots with believable data. This capability enables the creation of complete testing environments for satellite software that integrate a variety of simulated subsystems and sensors. For instance, Satellite Simulator can simulate orbital mechanics (e.g., position tracking, attitude adjustment) and electrical power systems (e.g., power generation, consumption, and distribution) alongside other subsystems.
We leveraged our framework to create honeypots of real-world CubeSat missions. Our results, backed by our survey of satellite operators, show that HoneySat’s simulation is highly realistic. Our framework simulates an entire satellite mission, provides realistic telemetry, and supports real telecommands.
Moreover, in a significant showcase of HoneySat’s practical relevance, a leading multinational space agency has integrated HoneySat into operational test environments to evaluate its effectiveness in securing mission-critical space systems.
Finally, we collaborated with an aerospace company to perform a hardware-in-the-loop experiment where HoneySat successfully communicated with an in-orbit, operational SmallSat.
@@
  • We introduce HoneySat, a novel, high-interaction, extensible honeypot framework for small satellites (Sec. IV).
  • We present the Satellite Simulator, a library that simulates the physical processes, sensors, and subsystems necessary for a realistic satellite honeypot (Sec. V).
  • We demonstrate HoneySat’s high level of realism through a hands-on survey where 10 experienced satellite operators interacted with HoneySat in real time (Sec. VI-C).
  • We show HoneySat’s extensibility features by integrating two real-world satellite flight software and integrating HoneySat into a hardware-in-the-loop operation allowing our honeypot to communicate with an in-orbit operational SmallSat (Sec. VI-E and Sec. VI-F).
II Background
This section lays out key background concepts that are relevant to satellite honeypots. For honeypots: a description of the different existing types (Sec. II-A), as well as the current state-of-the-art (Sec. II-B). For satellites: their operation (Sec. II-C), architecture (Sec. II-D), existing protocol ecosystems (Sec. II-E), and tactics, techniques, and procedures (TTPs) (Sec. II-F).
II-A Types of Honeypots
@@
Low-Interaction Honeypots. These honeypots offer minimal interaction, simulating real systems through scripts. They are easy to setup and maintain due to low resource consumption, and have a reduced risk of adversarial takeover. However, they provide limited interaction opportunities to adversaries, which limits the data they provide. Low-interaction honeypot examples include Conpot [30] and Honeyd [31].
High-Interaction Honeypots. These honeypots offer extensive interaction opportunities via emulation or advanced simulations [31]. Their main advantage is providing adversaries with almost limitless interactions, providing extensive interaction data. However, they pose a high risk of adversarial takeover as adversaries have more opportunities to hijack the honeypot [32]. High-interaction honeypot examples include Cowrie [33] and HoneyPLC [23].
II-B Honeypot’s State of the Art
The literature on honeypots includes hundreds of implementations that simulate a diverse set of systems [34, 35]. From implementations that simulate a host’s TCP/IP stack, such as Honeyd [31], to modern approaches that integrate social media applications, such as HoneyTweet [24]. In the absence of a satellite honeypot, we examine the approaches most related to satellites: Industrial Control Systems (ICS) and Unmanned Aerial Vehicles (UAV), summarized in Table I.
Satellite systems like ICS must be aware of some physical process, e.g., the sun’s position, via sensors to acquire data about the physical world. Several ICS honeypots have simulated these physical processes. For example, ICSnet [36] and HoneyICS [37] simulate components such as programmable logic controllers (PLCs) and actuators such as valves.
HoneyDrone [38] is a UAV honeypot that integrates different simulations, e.g., Ardupilot, to recreate multiple attack scenarios. Although UAV honeypots share some similarities with satellites [39], satellite honeypots require additional physical simulations and involve a more complex operation environment, which we discuss in Sec. II-C.
TABLE I: Comparison of Existing Honeypots and HoneySat.
Keys: = Supported; $\times$ = Not Supported.
@@ II-B Honeypot’s State of the Art
IV-D,
IV-C
IV-C,
V-B1
IV-D,
V-A
VI-E
VI-F
II-C Anatomy of a Satellite Mission
We now describe the components of a satellite mission. Due to satellite missions’ complexity and diversity, we explain each component and match it with one of the numbers in Fig. 1. Every satellite mission includes the ground segment from which satellite operators control the satellite and the space segment which includes the satellite itself.
Ground Segment 1. The Ground Segment (GS) covers the terrestrial infrastructure required for a successful satellite operation. It consists of a ground station, responsible for exchanging data with the spacecraft, the computational and network infrastructure required for communication, and the systems to operate the satellites, e.g., servers [12]. The GS includes the Ground Segment Software (GSS) that helps operators schedule and send commands and visualize data.
Telecommands (TC) 3 and Telemetry (TM) 4. The basic data flow between the space and ground segments are TC and TM [41]. TM is the data the satellite sends to the ground station which may contain the satellite’s status or payload data [41]. TCs are used to operate the satellite and are transmitted and encapsulated in a space protocol (see Sec. II-E). The design and implementation of TCs varies depending on the satellite mission. From a security perspective, TCs are particularly important as an attacker that can send valid TCs to a satellite can fully take over the mission [12].
Orbital Pass 5. A satellite and its ground station can communicate only during an orbital pass. An orbital pass, or more commonly a pass, is when the satellite rises above a ground station’s horizon and becomes available for communication. A pass’s duration and timing depend on the satellite’s orbit characteristics and any obstructing objects between the satellite and the ground station, e.g., mountains [42]. In some cases, such as geostationary orbit (GEO), the satellite can be above the horizon constantly. The occurrence of passes can be calculated using the spacecraft’s orbital elements in the two-line element (TLE) data [43].
Satellite Mission Operations 6. Satellite mission operations vary widely depending on the owning organization’s budget, and technology [5]. Nevertheless, they share some commonalities, which we now describe.
A mission’s operation involves a team of operators that use ground segment software (GSS) to operate a satellite(s) and ensure the mission’s success [44]. Operations are carried out in a Mission Operations Center (MOC), where operators sit at their workstations to manage TCs sent to the satellite(s). Operators may also remotely operate satellites 7 by connecting to the ground segment using tools such as VNC (Virtual Network Computing) [11].
Satellite operations include two main activities: satellite tracking and TC generation and scheduling. Satellite tracking calculates the satellite’s position in orbit and controls the ground station’s tracking antenna to establish communication between the ground and space segments. TC generation and scheduling crafts commands to be sent to the satellite to perform different functions, e.g., download payload data.
@@ II-C Anatomy of a Satellite Mission
Ground (Segment) Software (GSS). GSS allows operators to carry out the mission’s routine operations. GSS are very diverse. Some satellite missions develop their own GSS while others use open-source [45] or proprietary GSS [46]. There are two main types of GSS: Mission Control Software (MCS) and Ground Station Control Software (GSCS).
@@
Mission control software 8 manages TCs and scripts to be sent to the satellite and display TM. For example, ESA’s SCOS-2000 is an MCS that provides generic functionality that can be customized for a specific mission operation [47]. Some missions develop their own MCS. For example, the SUCHAI mission developed their own MCS application [45].
@@
Ground station control software (GSCS) 9 helps satellite operators track and visualize the satellite’s orbit and provide information about each satellite’s pass. For example, Gpredict is a popular open-source GSCS that performs real-time satellite tracking and orbit prediction [48].
II-D Satellite Architecture
@@
Satellite architectures are varied and complex [49]; however, here we describe the most common terminology depicted in Fig. 1’s space segment 2.
@@
When referring to satellite architecture there is a distinction between the platform that facilitates the successful operation of the satellite’s activities and the payload. The platform provides the possibility to run a payload that fulfills the purpose of these missions. Payloads are different depending on the satellite’s mission and can range from instruments to measure physical properties to communications systems providing wireless connectivity.
@@
Platform: The platform is composed of custom-designed or off-the-shelf subsystems necessary for critical satellite operations. These include the Attitude Determination and Control System (ADCS) to maintain the satellite’s orientation (i.e. attitude) and position to keep the satellite pointed towards antennas or solar panels illuminated; the Electrical Power Subsystem (EPS) for managing power generation and distribution; the Communication Subsystem (COMM) and the Command and Data Handling (C&DH) subsystems to facilitate communications for receiving TC and sending TM and controlling the satellite operations.
@@
These subsystems are controlled via TCs sent to the satellite. TCs may be interpreted by a central C&DH System or merely forwarded to the recipient subsystem for further processing [50]. This managerial duty is handled by the Flight Software (FS) running on an embedded system. Some examples include NASA’s Core Flight System (cFS) [51] and the University of Chile’s SUCHAI FS [52]. Attackers aim to gain the ability to send TCs to the flight software as this may grant control over all spacecraft subsystems.
@@
Payload. The payload is the equipment that a satellite employs to fulfill its mission. Due to satellites’ varied missions, payloads are heavily customized [53]. For example, if a satellite’s mission is remote sensing, its payload may include an infrared camera [54].
II-E Small Satellite Protocol Ecosystems